Project

General

Profile

Joint MICE Experiment Management Office and MICE International Project Office

24th July 2018; 15:30 BST

Phone details

Agenda

  1. Introduction: KL
  2. Notes of previous meeting and status of actions: All
    • KL: will migrate EB and CB mailing lists to jiscmail
    • CR: will migrate analysis mailing list to jiscmail
    • DR: will migrate software mailing list to jiscmail
  3. Project Manager's report: CW
  4. Software and computing status: DR
  5. Paper status and plans: CR
  6. AOB

Notes

Present: KL, ABr, DR, JB,

  1. Introduction: KL
    • Routine meeting to continue to address issues related to analysis.
  2. Notes of previous meeting and status of actions: All
    • Notes accepted. Status of actions:
    • KL: will migrate EB and CB mailing lists to jiscmail
      • Stands.
    • CR: will migrate analysis mailing list to jiscmail
      • Stands.
    • DR: will migrate software mailing list to jiscmail
      • Stands.
  3. Project Manager's report: CW
    • No report as CW did not attend. US equipment has been collected. Status of European shipments unknown.
  4. Software and computing status: DR
    • Principal issue is to sort out the TOF. V.Pec has checked the most recent runs. Now cross-validating the calibration etc. of previous runs. Also, plan to loosen time-correlation cut in the reconstruction. Now keep spacepoints if the times in the two slabs agree to within 3ns.
    • Monte Carlo production with new MAUS version has moved to the GRID. Some issues are being sorted out before general-purpose MC production was launched. Data processing with new MAUS is also underway.
    • When TOF calibration is sorted out data will be reprocessed. Resources are limitted as VP is now also working on DUNE. Data/MC comparison for TOF needs to be expedited.
    • C.Hunt has developed a new calibration scheme for the tracker. That means that a new iteration for the tracker calibration will be produced. The data will have to be reprocessed.
    • The density of the tracker stations is a data-card input to MAUS. A tuning of the relevant parameter will have to be under taken.
    • The cuts data structure is now in MAUS but it is not yet in production in MAUS. The issue is that the handling in the processing is not straightforward. The cuts will have to be defined and then the implementation finalised.
    • Need to finalise the magnetic-field scaling for use in the various cooling-channel settings. This implies a systematic procedure to deal with each of the settings.
    • The use of the TOF to seed a re-fit of low-pt tracks needs to be revisitted.
    • Geometry handover with P.Franchini is in process. Processing scripts are in the repository. Documentation must be provided.
    • The MAUS paper has received many comments from D.Kaplan. One particular comment was that the computer jargon needs to be explained. DR expects to issue a new version by the end of this week.
  5. Paper status and plans: CR
    • CR unavailable. Points noted in the discussion:
      • C.Hunt with V.Blackmore has found a means to re-weight the MC at the handoff point between G4BeamLine and MAUS. This seems to improve the agreement. More detailed studies are under way.
      • CR now on has a seeding algorithm for the low-pt tracks downstream.
      • Agree to consider PRAB or EU.J.Phys.C for the emittance-measurement paper.
  6. AOB
    • None.

Summary of actions:

  • KL: will migrate EB and CB mailing lists to jiscmail.
  • CR: will migrate analysis mailing list to jiscmail.
  • DR: will migrate software mailing list to jiscmail.

Back to MEMO:

Back

Back to MIPO:

Back

Documents